Asset A pays $100.00 in dividends each year, which are taxed at 25%. Other than this annual payment, asset A does not increase in value each year. Asset B pays no dividends, but you must pay a 25% tax on its total value when you sell it. The market interest rate is 10%.

a) Suppose you purchase asset A intending to keep it for 10 years (receiving 10 total dividend payments), at which time you will sell it for the same price you purchased it for. What is the present value of the returns from that purchase?

b) Suppose you receive asset B for free and plan to keep it for 10 years and then sell it, at which point its value will have risen to $1,000.00. What is the present value of the return from that asset?
Answer #1

a. the dividends after tax = 100*(1 - 0.25) = 75

FV = 0, N = 10, PMT = 75, rate = 10%

use PV function in Excel

present value of A = 460.84

b) Value after 10 years = 1000*(1 - 0.25) = 750

present value = 750/1.10^10 = 289.16